OpenStudy (anonymous):

If g(x) = 3x + 5, what is g^–1(x)?

OpenStudy (helder_edwin):

if u mean \[ \large g^{-1}(x) \] then just solve for \(x\) the equation \[ \large y=g(x)=3x+5 \]

OpenStudy (anonymous):

YES

OpenStudy (helder_edwin):

solve the equation and tell what u got

OpenStudy (anonymous):

X=-1 2/3

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

u have \[ \large y=3x+5 \] first substract 5 from both sides

OpenStudy (anonymous):

x= -5+y/3

OpenStudy (helder_edwin):

u r missing some parenthesis!!!!!!

OpenStudy (anonymous):

where

OpenStudy (anonymous):

g-1= ( g(x) - 5 )/3

OpenStudy (helder_edwin):

do u see the difference between \[ \large x=-5+\frac{y}{3} \] (this is what u wrote) and \[ \large x=\frac{-5+y}{3}=(-5+y)/3 \]

OpenStudy (helder_edwin):

which one is correct??

OpenStudy (anonymous):

the second one

OpenStudy (helder_edwin):

yes!! do u see how important parentheses are??? they can change completely an expression!!!!!

OpenStudy (helder_edwin):

so \[ \large g^{-1}(y)=\frac{y-5}{3} \]